# Bootstrap Dynamic-Aware 3D Visual Representation for Scalable Robot Learning

### v2

@@ 4.3 Variance-Regularized Representation Matching
Here, $\mathcal{L}_{\text{inv}}$ enforces alignment between student and teacher features, $\mathcal{L}_{\text{var}}$ maintains sufficient variance along each dimension to avoid collapse, and $\mathcal{L}_{\text{cov}}$ reduces cross-channel correlations for disentangled representations. Together, they stabilize latent dynamics and improve the encoder’s ability to capture coherent visual structures.
@@ 4.4 Pre-training Details
Table 1: Quantitative comparison on simulation benchmarks. Type denotes the pretraining category: 2D-LT = large-scale 2D pretraining; 3D-LT = large-scale 3D pretraining; 3D-Pre-S = static 3D pretraining without dynamic modeling; 3D-Pre-D = dynamic-aware 3D pretraining with temporal modeling; Input Type: RGB = image input; PC = point cloud input. Conclusion: AFRO achieves the highest success rate across all task categories, demonstrating that our pre-training method can substantially enhance the performance in robot manipulation.
@@
CLIP [clip_ref] 2D-LT RGB 29.3 14.0 5.0 55.3 24.9 ($\downarrow$ 51.1) 61.0 84.0 72.5 ($\downarrow$ 10.5)
DINOv2 [dinov2_ref] 2D-LT RGB 34.0 14.0 5.0 55.3 25.9 ($\downarrow$ 50.1) 76.0 84.0 80.0 ($\downarrow$ 3.0)
PointNet [pointnet_ref] 3D-LT PC 71.3 49.7 23.0 55.3 51.7 ($\downarrow$ 24.3) 80.0 72.0 76.0 ($\downarrow$ 7.0)
PointMAE [pointmae_ref] 3D-Pre-S PC 77.3 61.3 42.0 70.0 63.9 ($\downarrow$ 12.1) 64.0 76.0 70.0 ($\downarrow$ 13.0)
PointDif [pointdiff_ref] 3D-Pre-S PC 83.3 58.0 41.0 58.7 61.1 ($\downarrow$ 14.9) 76.0 78.0 77.0 ($\downarrow$ 6.0)
DynaMo [dynamo_ref] 2D-Pre-D RGB 42.0 21.7 14.0 34.0 27.6 ($\downarrow$ 48.4) 76.0 68.0 72.0 ($\downarrow$ 11.0)
DynaMo-3D [dynamo_ref] 3D-Pre-D PC 84.7 55.3 41.0 80.0 64.9 ($\downarrow$ 11.1) 73.0 76.0 74.5 ($\downarrow$ 8.5)
FVP [fvp_ref] 3D-Pre-D PC 80.0 44.3 34.0 62.0 54.3 ($\downarrow$ 21.7) 66.0 76.0 71.0 ($\downarrow$ 12.0)
DP3 [dp3_ref] 3D-Pol PC 82.7 65.0 49.0 80.0 69.7 ($\downarrow$ 6.3) 70.0 80.0 75.0 ($\downarrow$ 8.0)
@@ 5 Simulation Experiment
Adroit [adroit_ref] is a high-fidelity multi-DoF anthropomorphic hand benchmark in MuJoCo focused on fine-grained control. We evaluate two tasks, Door and Pen, both of which require precise manipulation and temporal coordination.
MetaWorld [metaworld_ref] provides a diverse set of manipulation tasks using a 7-DoF Sawyer arm in MuJoCo. We select 14 representative tasks (all details in Appendix).
In MetaWorld, each task includes 25 expert trajectories generated using scripted policies, while in Adroit each task contains 100 expert trajectories produced by VRL3 [vrl3_ref].
@@
To evaluate the effectiveness of AFRO, we compare it against four categories of baselines. (1) End-to-End Learning (DP3) [dp3_ref] jointly learns perception and control without visual pretraining. (2) Large-Scale Dataset Pretraining employs visual feature extractors such as CLIP [clip_ref], DINOv2 [dinov2_ref], and PointNet [pointnet_ref] trained on large external datasets to improve generalization before diffusion policy training. (3) Static-Only Pretraining baselines, including Point-MAE [pointmae_ref] and PointDif [pointdiff_ref], rely solely on static data, ignoring dynamic state relationships. (4) Dynamic-Aware Pretraining baselines, such as FVP [fvp_ref], DynaMo-2D [dynamo_ref], and DynaMo-3D, explicitly model temporal dynamics during visual pretraining, with DynaMo-3D being our 3D extension of the original DynaMo method.
@@ 5.4 Quantitative Results
As shown in Table 1, AFRO achieves state-of-the-art performance on both benchmarks under a unified training and evaluation protocol. Large-scale 2D models (CLIP, DINOv2) provide strong semantic priors but perform poorly on MetaWorld (24.9%, 25.9%). Among 3D methods, static approaches (PointMAE, PointDif) benefit from geometric reconstruction yet lack temporal reasoning, while dynamic-aware baselines such as DynaMo-3D can even underperform static models. In contrast, AFRO couples diffusion-based future-feature prediction with causally structured latent-action learning, reaching 76.0% mean success on MetaWorld (+6.3 over DP3) and 83.0% on Adroit. The especially large gains on challenging tasks indicate that our feature differencing and inverse-consistency supervision yield stable, manipulation-oriented visual representations that better capture the causal structure of state transitions.
Figure 4: Scaling across task domains. Comparison between in-domain (solid) and multi-domain (hatched) visual pretraining on four tasks. AFRO consistently benefits from multi-domain pretraining and even reaches $100\%$ success on Peg Unplug Side, whereas static and dynamic baselines show smaller or inconsistent gains.
@@ 5.5 Scalability Analysis
Figure 5: Scaling with data on Adroit. Success rate versus number of expert trajectories (10, 20, 50, 100, 200, 500) for Door and Pen.
Scalable Evaluation of Data. We vary the number of expert trajectories per task $\{10,20,50,100,200,500\}$ on Adroit-Door and Adroit-Pen (Fig. 5). AFRO improves smoothly and dominates across all data scales. On Door, it reaches 74% with only 50 samples and 90% at 500, surpassing DP3 and static pretraining baselines that plateau around 70–76%. On Pen, AFRO jumps from 60% to 81% when increasing samples from 20 to 50, and further to 98% at 500, while alternatives fluctuate or saturate early. This indicates that AFRO leverages additional data to learn richer latent dynamics, mapping unseen states to familiar transition patterns and enabling more effective policy decoding.
Scalable Evaluation of Domain. We evaluate how visual pretraining scales with task diversity. To this end, we conduct a multi-domain setting where the encoder is pretrained once on the union of tasks (Door+Pen for Adroit, all 14 tasks for MetaWorld) and then frozen while a diffusion policy is trained on the target task. As shown in Fig. 4, AFRO benefits substantially from multi-domain pretraining: average success rises from 68.0% to 82.8%, with Bin Picking improving from 20% to 58% and Peg Unplug Side reaching 100%, whereas other methods show only modest gains or even degrade. This indicates that AFRO better exploits heterogeneous trajectories to learn transferable 3D dynamics, organizing states by transition structure rather than superficial appearance.
Figure 6: t-SNE visualization of latent features on eight MetaWorld tasks. Colors denote tasks and marker shapes denote different demonstration sequences. AFRO (left) forms clearer task clusters than DynaMo-3D (right), indicating more discriminative and temporally coherent representations.
@@ 6 Real-World Experiment
Figure 7: Real-world evaluation setup and objects. (a) Franka Emika arm with a parallel gripper and the objects used across tasks. (b) Physical setup with a top-down RealSense L515 RGB-D camera. (c) Size references for the main objects (dimensions in cm).
@@ 6.1 Experimental Setup
We evaluate on a manipulation platform built around a 7-DoF Franka Emika arm with a parallel gripper and a top-down RealSense L515 depth camera; object layouts and camera/robot placement follow the real-world configuration in Fig. 7. We consider four representative manipulation tasks spanning non-prehensile and prehensile control, perception precision, and kinematic difficulty:
  • Block-to-Block Alignment: Using non-prehensile manipulation, the robot pushes a movable block to precisely align its edge with a reference block. (B2BA)
  • Bell Pressing: The robot localizes a physical bell and presses it accurately to trigger an audible click. (BP)
  • Fruit Pick-and-Place: The robot grasps a fruit from a randomized start pose and places it into a target basket. (FPP)
  • Cover Block: The robot picks up a cup from a source location, moves above a target block, and places the cup to fully cover the block. (CB)
For all tasks, shown in Fig. 8, object initializations are randomized within bounded perturbations to assess robustness. For each task, we collect 40 demonstrations, and we report success rates over 20 evaluation trials. We use a PointTransformer [pointtransformer] backbone as the visual feature encoder for all real-world experiments.
@@ 6.2 Real-World Task Performance
As shown in the blue block of Table 2, AFRO achieves the highest mean success rate of 0.70 across all four real-world tasks, outperforming DP3 (0.50), PointDif (0.53), FVP (0.34), and DynaMo-3D (0.48). The largest gains appear on Fruit Pick-and-Place and Cover Block, which involve wide-range workspace motions (e.g., lifting a fruit from the table into an elevated basket and transporting a cup over a target block). These results indicate that AFRO’s transition-centric 3D representations transfer robustly from simulation to a physical Franka setup, especially under large-magnitude spatial movements and real-world sensing noise.
@@ 6.3 Large-Scale Out-of-Domain Pretraining
We further evaluate scalability to large, out-of-domain data by pretraining on the RH20T [rh20t] real-world manipulation dataset. Specifically, we construct a Franka subset (RH20T_Franka_Simple) by reconstructing point clouds from camera intrinsics and depth, cropping to a fixed workspace, discarding the first 30 static frames, and sampling one frame every 20 from the first two scenes of over 140 tasks. On this corpus, we pretrain a shared PointTransformer encoder with AFRO, DynaMo-3D, PointDif, and FVP, then fine-tune it on four real-robot manipulation tasks with a diffusion action decoder. As shown in the red block of Table 2, AFRO benefits the most from large-scale out-of-domain pretraining, raising mean success from 70.0% to 84% and clearly outperforming all competing methods, indicating a stronger ability to exploit heterogeneous real-world demonstrations for transferable 3D manipulation dynamics.
@@ Appendix B Additional Related Work
Robotic manipulation aims to transform object states in the physical world, from basic grasping[zheng2024gaussiangrasper, dexsingrasp] to long-horizon, contact-rich [wu2024rttf] tasks. Early systems decomposed perception, grasp planning, and motion planning, while recent work learns end-to-end policies from rich sensory inputs such as RGB(-D) [chi2025diffusion], tactile [wu2024tars, zhou2025gentle, huang2025tactile], and force feedback [forcefb], often with additional safety constraints [Wang_2025_ICCV] or language conditioning. Vision–language–action models [kim2024openvla] further condition policies on natural-language goals but are still mostly built on 2D image encoders, which struggle with occlusions and precise 6-DoF reasoning. To overcome these limitations, recent approaches use 3D-aware policies that operate on depth, multi-view RGB-D, or point clouds, such as methods that infer volumetric features or explicit 3D action maps [gervet2023act3d, rvt, shridhar2023perceiver]. Our work is complementary: instead of introducing another policy head, AFRO provides a 3D dynamics-aware representation that can be plugged into existing diffusion-based manipulation policies with minimal changes.
@@ Appendix E Further Details on RH20T
RH20T[rh20t] (Robot–Human demonstration in 20TB) is a large-scale, real-world robotic manipulation dataset designed to support one-shot imitation learning and general skill learning across diverse, contact-rich tasks. The dataset contains over 110,000110{,}000 robot manipulation sequences paired with an equal number of human demonstration videos, resulting in more than $50$ million image frames in total and about $20$ TB of data. It demonstrates over 140 different tasks, including a wide variety of operational tasks. Each robot sequence is recorded in the real world and includes synchronized multi-modal information such as RGB and depth images, force/torque measurements, audio, and low-level robot state and action signals, together with a corresponding human demonstration video and a language description for the same skill. We generate a point cloud for each frame by back-projecting the depth map into 3D space using the corresponding camera intrinsics. We then apply farthest point sampling (FPS) to downsample the raw point cloud to 1024 points. Empirically, we find that 1024 points provide sufficient geometric coverage to capture the object contours while at the same time substantially reducing the computational cost of the PointTransformer encoder.
@@ Appendix F Limitations and Future Work
These observations suggest several directions for future work. (1) Scaling and semantic grounding. We plan to combine AFRO with strong semantic backbones such as VGGT [wang2025vggt] or DINOv2 style models, either by distilling their features into the 3D encoder or by joint multi-task pre-training. This would aim to obtain representations that are simultaneously dynamics-aware and semantically rich. (2) Larger-scale pre-training. A natural extension is to train AFRO on substantially larger and more diverse 3D robot datasets, including multi-embodiment, multi-task, and web-scale synthetic data, to approach the scale of general-purpose visual foundation models for robotics. (3) Multi-view 3D dynamics. Incorporating multi-view inputs and explicitly modeling cross-view consistency of dynamics (e.g., by fusing trajectories from several cameras) could make the representation more robust to occlusion and viewpoint changes and better suited to mobile manipulation. (4) Physically grounded latent actions. We are interested in tying latent actions to physically interpretable quantities, for example by supervising them with estimated optical flow, 3D scene flow, or end-effector trajectories, so that the latent space captures meaningful motion primitives that can be reused across tasks and robots. (5) Collapse-free objectives beyond VICReg. Finally, we aim to explore alternative self-supervised objectives—such as JEPA-style prediction, information-theoretic regularization, or architectural constraints—that maintain feature diversity without relying on explicit variance penalties, potentially simplifying training and further improving robustness of the learned 3D representations.

### v3

@@ 4.3 VICReg Matching
Here, $\mathcal{L}_{\text{inv}}$ aligns student and teacher features, $\mathcal{L}_{\text{var}}$ maintains variance, and $\mathcal{L}_{\text{cov}}$ reduces cross-channel correlations, stabilizing latent dynamics and improving representations.
@@ 4.4 Pre-training Details
Table 1: Comparison on simulation benchmarks. Type denotes the pretraining category: 2D-LT = large-scale 2D pretraining; 3D-LT = large-scale 3D pretraining; 3D-Pre-S = static 3D pretraining without dynamic modeling; 3D-Pre-D = dynamic-aware 3D pretraining with temporal modeling; Input Type: RGB = image input; PC = point cloud input. Conclusion: AFRO achieves the highest success rate across task categories, demonstrating our pre-training method can substantially enhance performance in robot manipulation.
@@
CLIP [35] 2D-LT RGB 29.3 14.0 5.0 55.3 24.9 ($\downarrow$ 51.1) 61.0 84.0 72.5 ($\downarrow$ 10.5)
DINOv2 [29] 2D-LT RGB 34.0 14.0 5.0 55.3 25.9 ($\downarrow$ 50.1) 76.0 84.0 80.0 ($\downarrow$ 3.0)
PointNet [33] 3D-LT PC 71.3 49.7 23.0 55.3 51.7 ($\downarrow$ 24.3) 80.0 72.0 76.0 ($\downarrow$ 7.0)
PointMAE [31] 3D-Pre-S PC 77.3 61.3 42.0 70.0 63.9 ($\downarrow$ 12.1) 64.0 76.0 70.0 ($\downarrow$ 13.0)
PointDif [62] 3D-Pre-S PC 83.3 58.0 41.0 58.7 61.1 ($\downarrow$ 14.9) 76.0 78.0 77.0 ($\downarrow$ 6.0)
DynaMo [30] 2D-Pre-D RGB 42.0 21.7 14.0 34.0 27.6 ($\downarrow$ 48.4) 76.0 68.0 72.0 ($\downarrow$ 11.0)
DynaMo-3D [30] 3D-Pre-D PC 84.7 55.3 41.0 80.0 64.9 ($\downarrow$ 11.1) 73.0 76.0 74.5 ($\downarrow$ 8.5)
FVP [15] 3D-Pre-D PC 80.0 44.3 34.0 62.0 54.3 ($\downarrow$ 21.7) 66.0 76.0 71.0 ($\downarrow$ 12.0)
DP3 [59] 3D-Pol PC 82.7 65.0 49.0 80.0 69.7 ($\downarrow$ 6.3) 70.0 80.0 75.0 ($\downarrow$ 8.0)
@@ 5 Simulation Experiment
Adroit [38] is a multi-DoF anthropomorphic hand benchmark in MuJoCo focused on fine-grained control. We evaluate two tasks, Door and Pen, both of which require precise manipulation and temporal coordination.
MetaWorld [57] provides a diverse set of manipulation tasks using a 7-DoF Sawyer arm in MuJoCo. We select 14 representative tasks (all details in Appendix).
In MetaWorld, each task includes 25 expert trajectories generated using scripted policies, while in Adroit each task contains 100 expert trajectories produced by VRL3 [42].
@@
To evaluate the effectiveness of AFRO, we compare it against four categories of baselines. (1) End-to-End Learning (DP3) [59] jointly learns perception and control without visual pretraining. (2) Large-Scale Dataset Pretraining employs visual feature extractors such as CLIP [35], DINOv2 [29], and PointNet [33] trained on large external datasets to improve generalization before diffusion policy training. (3) Static-Only Pretraining baselines, including Point-MAE [31] and PointDif [62], rely solely on static data, ignoring dynamic state relationships. (4) Dynamic-Aware Pretraining baselines, such as FVP [15], DynaMo-2D [30], and DynaMo-3D, explicitly model temporal dynamics during visual pretraining, with DynaMo-3D being our 3D extension of the original DynaMo method.
@@ 5.4 Quantitative Results
As shown in Table 1, AFRO achieves the best performance on both benchmarks under a unified protocol. 2D models (CLIP, DINOv2) provide semantic priors but perform poorly on MetaWorld (24.9%, 25.9%). Among 3D methods, static approaches (PointMAE, PointDif) benefit from geometric reconstruction but lack temporal reasoning, while dynamic baselines such as DynaMo-3D underperform static models. In contrast, AFRO combines diffusion-based future prediction with latent-action learning, reaching 76.0% success on MetaWorld (+6.3 over DP3) and 83.0% on Adroit. These gains suggest feature differencing and inverse-consistency supervision yield stable representations capturing causal state transitions.
Figure 4: Scaling across task domains. Comparison of in-domain (solid) and multi-domain (hatched) visual pretraining. AFRO benefits from multi-domain data and reaches $100\%$ success on Peg Unplug Side, while baselines show smaller or inconsistent gains.
@@ 5.5 Scalability Analysis
Figure 5: Scaling with data. Success rate versus number of expert trajectories (10, 20, 50, 100, 200, 500) for Door and Pen.
Scalable Evaluation of Domain. We study how pretraining scales with task diversity. The encoder is pretrained once on the union of tasks (Door+Pen for Adroit, 14 tasks for MetaWorld) and then frozen while a diffusion policy is trained for each task. As shown in Fig. 4, AFRO benefits from multi-domain pretraining: mean success rises from 68.0% to 82.8%, with Bin Picking improving from 20% to 58% and Peg Unplug Side reaching 100%. Other methods show smaller gains or degradation. This indicates AFRO learns transferable 3D dynamics by organizing states through transition structure instead of appearance. Overall, AFRO scales favorably with increasing task diversity and heterogeneous training domains.
Figure 6: t-SNE of latent features on eight MetaWorld tasks. Colors denote tasks and marker shapes denote sequences. AFRO (left) forms clearer clusters than DynaMo-3D (right), indicating more discriminative and temporally coherent representations.
Scalable Evaluation of Data. We vary expert trajectories per task $\{10,20,50,100,200,500\}$ on Adroit-Door and Adroit-Pen (Fig. 5). AFRO leads across all scales. On Door, it reaches 74% with 50 samples and 90% at 500, exceeding DP3 and static baselines that plateau around 70–76%. On Pen, AFRO rises from 60% to 81% when samples increase from 20 to 50, and to 98% at 500, while alternatives fluctuate or saturate early. This suggests AFRO leverages additional data to learn richer latent dynamics for more effective policy decoding. Overall, AFRO exhibits consistently favorable scaling behavior as data increases.
@@ 6 Real-World Experiment
Figure 7: Real-world evaluation setup and objects.
@@ 6.1 Experimental Setup
We evaluate on a platform with a 7-DoF Franka Emika arm, parallel gripper, and a top-down RealSense L515 depth camera; object layouts and camera/robot placement follow Fig. 7. We consider four manipulation tasks covering non-prehensile and prehensile control, perception precision, and kinematic difficulty:
  • Block-to-Block Alignment: The robot pushes a movable block to align it with a reference block. (B2BA)
  • Bell Pressing: The robot localizes a bell and presses it to trigger a click. (BP)
  • Fruit Pick-and-Place: The robot grasps a fruit from a random pose and places it into a target basket. (FPP)
  • Cover Block: The robot picks up a cup, moves above a target block, and places the cup to cover it. (CB)
For all tasks (Fig. 8), initial object poses are randomized within bounded perturbations to test robustness. For each task we collect 40 demonstrations and report success rates over 20 trials. A PointTransformer [60] backbone serves as the visual encoder for all real-world experiments.
@@ 6.2 Real-World Task Performance
As shown in the blue block of Table 2, AFRO achieves the highest mean success rate of 0.70 across four tasks, outperforming DP3 (0.50), PointDif (0.53), FVP (0.34), and DynaMo-3D (0.48). The largest gains occur on Fruit Pick-and-Place and Cover Block, which require wide workspace motions (e.g., lifting fruit into an elevated basket or moving a cup over a target block). These results indicate that AFRO’s transition-centric 3D representations transfer reliably from simulation to a Franka setup under large spatial motions and real-world sensing noise.
@@ 6.3 Large-Scale Out-of-Domain Pretraining
We further evaluate scalability to large, out-of-domain data by pretraining on the RH20T [11] real-world manipulation dataset. Specifically, we construct a Franka subset (RH20T_Franka_Simple) by reconstructing point clouds from camera intrinsics and depth, cropping to a fixed workspace, discarding the first 30 static frames, and sampling one frame every 20 from the first two scenes of over 140 tasks. On this corpus, we pretrain a shared PointTransformer encoder with AFRO, DynaMo-3D, PointDif, and FVP, then fine-tune it on four real-robot manipulation tasks with a diffusion action decoder. As shown in the red block of Table 2, AFRO consistently benefits the most from large-scale out-of-domain pretraining, raising mean success from 70.0% to 84% and clearly outperforming all competing methods, indicating a stronger ability to effectively exploit heterogeneous real-world demonstrations for transferable 3D manipulation dynamics.
@@ Appendix B Additional Related Work
Robotic manipulation aims to transform object states in the physical world, from basic grasping[63, 51] to long-horizon, contact-rich [47] tasks. Early systems decomposed perception, grasp planning, and motion planning, while recent work learns end-to-end policies from rich sensory inputs such as RGB(-D) [7], tactile [48, 64, 16], and force feedback [45], often with additional safety constraints [44] or language conditioning. Vision–language–action models [22] further condition policies on natural-language goals but are still mostly built on 2D image encoders, which struggle with occlusions and precise 6-DoF reasoning. To overcome these limitations, recent approaches use 3D-aware policies that operate on depth, multi-view RGB-D, or point clouds, such as methods that infer volumetric features or explicit 3D action maps [13, 14, 39]. Our work is complementary: instead of introducing another policy head, AFRO provides a 3D dynamics-aware representation that can be plugged into existing diffusion-based manipulation policies with minimal changes.
@@ Appendix E Further Details on RH20T
RH20T[11] (Robot–Human demonstration in 20TB) is a large-scale, real-world robotic manipulation dataset designed to support one-shot imitation learning and general skill learning across diverse, contact-rich tasks. The dataset contains over 110,000110{,}000 robot manipulation sequences paired with an equal number of human demonstration videos, resulting in more than $50$ million image frames in total and about $20$ TB of data. It demonstrates over 140 different tasks, including a wide variety of operational tasks. Each robot sequence is recorded in the real world and includes synchronized multi-modal information such as RGB and depth images, force/torque measurements, audio, and low-level robot state and action signals, together with a corresponding human demonstration video and a language description for the same skill. We generate a point cloud for each frame by back-projecting the depth map into 3D space using the corresponding camera intrinsics. We then apply farthest point sampling (FPS) to downsample the raw point cloud to 1024 points. Empirically, we find that 1024 points provide sufficient geometric coverage to capture the object contours while at the same time substantially reducing the computational cost of the PointTransformer encoder.
@@ Appendix F Limitations and Future Work
These observations suggest several directions for future work. (1) Scaling and semantic grounding. We plan to combine AFRO with strong semantic backbones such as VGGT [43] or DINOv2 style models, either by distilling their features into the 3D encoder or by joint multi-task pre-training. This would aim to obtain representations that are simultaneously dynamics-aware and semantically rich. (2) Larger-scale pre-training. A natural extension is to train AFRO on substantially larger and more diverse 3D robot datasets, including multi-embodiment, multi-task, and web-scale synthetic data, to approach the scale of general-purpose visual foundation models for robotics. (3) Multi-view 3D dynamics. Incorporating multi-view inputs and explicitly modeling cross-view consistency of dynamics (e.g., by fusing trajectories from several cameras) could make the representation more robust to occlusion and viewpoint changes and better suited to mobile manipulation. (4) Physically grounded latent actions. We are interested in tying latent actions to physically interpretable quantities, for example by supervising them with estimated optical flow, 3D scene flow, or end-effector trajectories, so that the latent space captures meaningful motion primitives that can be reused across tasks and robots. (5) Collapse-free objectives beyond VICReg. Finally, we aim to explore alternative self-supervised objectives—such as JEPA-style prediction, information-theoretic regularization, or architectural constraints—that maintain feature diversity without relying on explicit variance penalties, potentially simplifying training and further improving robustness of the learned 3D representations.